# Multidimensional arrow of time

### v1

@@ Abstract
This paper studies the effect of extra dimensions on the arrow of time within the framework of $f(R)$ gravity. We demonstrate that the observed irreversibility of physical processes can be explained by the monotonic growth of the extra-dimensional space. Unlike traditional cosmological approaches, our model does not link the arrow of time to the entropy of matter or radiation; rather, it identifies it with the Bekenstein-Hawking-Wald entropy of the geometric background. We establish a formal relation between the volume of the multidimensional manifold and the statistical weights of its geometric states. This leads to a fundamental relationship where the flow of time is intrinsically linked to the growth of multidimensional entropy.
A key consequence of our framework is that the arrow of time remains a persistent feature for a 4D observer situated on a brane, even in the complete absence of matter or radiation. This directionality is driven by the dominant entropy production in the higher-dimensional bulk, which dominates local statistical fluctuations and ensures a stable causal direction.
@@ 2.1 Causality
The idea of a time-asymmetric layer of physical laws beneath the general relativity and quantum theory, was proposed by Roger Penrose in 1979 [11]. This idea was supported by Marina Cortes and Lee Smolin [7] where they introduced the existence of a large number of discrete states $S=I;J;K;L;$... - forming energetic causal sets endowed by an evolution rule. Each state, $I$, has a unique successor $J$. Similar idea is developed in the framework of causal sets [6, 13] where discrete states $A,B,C,\ldots$, forming a causal set $\mathcal{A}$. The order $A\prec B\prec C\prec\ldots$ denotes a ”precedes” causal relation. This idea is widely discussed in the literature, see e.g. [2]. The physical essence of ”state”  can be endowed by the energy and momentum [7], the discrete structure of the 3D space [14], or the entanglement entropy [15]. The importance of entropy in this concern was studied in [16]
@@ 3 The multidimensional arrow of time
Accidental metric fluctuations destroy any classical structures of the scale $l>1/H$. Therefore, scale of branes is limited by the $D$-dim horizon. It is clearly visible in the static de Sitter coordinates. In $D$-dim case this metric reads
$ds^{2}=\left(1-H^{2}\rho^{2}\right)dT^{2}-\frac{d\rho^{2}}{1-H^{2}\rho^{2}}-\rho^{2}d\Omega^{2}_{D-2},$
tending to the D-dim Minkowski metric at $\rho\ll 1/H$ where the quantum fluctuations are suppressed. If $l\ll 1/H$ i.e. deeply under the horizon, classical structures could be stable. Some of them can be found in [41].
@@
The solutions to classical equations (9) with the same action (8) describing the branes (18) and the expanding extra dimensions (11) have been studied separately, but with the same function of $f(R)$ gravity and the dimensionality of extra space, in [30] and [41]. The similar research can be found in [44], [45]. This approach is sufficient for our purpose, albeit at the cost of losing a smooth transition between the two regimes. The analogy follows the standard practice in cosmology: if our aim is to study the structure of a galaxy, we neglect processes like spatial expansion near the galaxy and solve the equations using the Minkowski metric. To study cosmological expansion, we consider the FRW metric as the solution to the Friedmann equations far from any local object. Matching these solutions at the boundary is of special interest and is not usually applied.
A reasonable physical analogy is provided by the Schwarzschild–de Sitter spacetime, described by the well-known metric:
$ds^{2}=f(\rho)dT^{2}-f(\rho)^{-1}d\rho^{2}-\rho^{2}d\Omega_{2}^{2},\quad f(\rho)=1-\frac{2M}{\rho}-H^{2}\rho^{2}.$ (19)
In this single coordinate system, two distinct spatial regimes coexist. While the positive cosmological constant $\Lambda$ drives accelerated expansion on cosmological scales, its influence is negligible in the strong-field region near the black hole. A clear criterion for this separation arises from comparing the gravitational potentials in the metric function $f(\rho)$. The term $2M/\rho$ represents the local mass potential, while $H^{2}\rho^{2}$ represents the repulsive potential associated with $\Lambda$. The transition radius $\rho_{*}$, where these influences are comparable, is found by equating their magnitudes:
$\frac{2M}{\rho_{*}}\sim H^{2}\rho_{*}^{2}\quad\Rightarrow\quad\rho_{*}\sim\left(\frac{2M}{H^{2}}\right)^{1/3}.$ (20)
@@
For radii $\rho\ll\rho_{*}$, the $H^{2}\rho^{2}$ term is negligible, defining a non-expanding region. In our model, the brane plays a role similar to this region: the brane remains a stable, static entity deeply ”under the horizon,” while the global bulk continues its expansion, thereby maintaining the established arrow of time.
@@
The Past Hypothesis postulates that the initial entropy of the Universe was extremely small compared to the entropy within the cosmological horizon today, [46]. Let us estimate this ratio within the framework of this work, using equation (17):
@@
Here, $S^{(D)}(0)$ denotes the Wald entropy contained within the cosmological horizon at the moment our brane was created. This initial value requires no fine-tuning within the framework of our model. $S^{(D)}(t)$ represents the Wald entropy of the same comoving volume at the present cosmic time, $t$.
@@
The entropy $S_{D}$, defined in Eq. (15), exhibits permanent growth in the multidimensional space. This constant flow of entropy dominates over any local entropy changes on the brane and supplies the arrow of time indefinitely.
@@ 5 Conclusion
This paper investigates the role of extra dimensions in the emergence of the arrow of time. We have shown that an initial volume, defined in a space with a $D$-dimensional de Sitter metric, grows continuously, generating disconnected volumes bounded by horizons. Each such volume is endowed with Bekenstein-Hawking-Wald entropy. Consequently, the total entropy of the initial volume increases exponentially along the temporal coordinate.
@@
Observers are located on a static brane with a scale significantly smaller than the de Sitter horizon. Since matter is, by definition, confined to the brane, the bulk expansion remains unobservable to a local observer. However, the entropy growth of the extra dimensions provides a rigorous basis for the arrow of time. To this end, following Penrose’s idea, we introduce a Postulate of Causality related to the entropy of physical states. We demonstrate that the arrow of time arises from the fact that the multidimensional volume is a monotonic function of the temporal coordinate. Conversely, the temporal coordinate $t$ is a monotonic function of entropy (16); therefore, applying the Postulate of Causality formally establishes the arrow of time.
@@ Acknowledgment
The author is grateful to V. Berezin for his interest in the research. The work was partly funded by the Ministry of Science and Higher Education of the Russian Federation, Project ”Studying physical phenomena in the micro- and macro-world to develop future technologies”  FSWU-2026-0010 and the Kazan Federal University Strategic Academic Leadership Program.

### v2

@@ Abstract
This paper investigates the influence of extra dimensions on the nature of the arrow of time. We demonstrate that the observed arrow of time can be explained by the monotonic growth of the multidimensional manifold’s volume. Unlike traditional cosmological approaches based on the entropy of matter or radiation, our model identifies the primary temporal direction with the Bekenstein-Hawking-Wald entropy of the geometric background.
By establishing a formal relation between the statistical weight of the multidimensional manifold and the multiplicity of causally disconnected regions, we reveal that time’s directionality is driven by dominant entropy production in the higher-dimensional bulk. A key consequence of this approach is that the arrow of time remains a persistent feature for a 4D observer even in the vacuum limit. This global geometric evolution suppresses local statistical fluctuations and ensures a robust and stable entropy flow throughout the manifold.
@@ 2.1 Causality
The concept of a time-asymmetric layer of physical laws underlying general relativity and quantum theory was proposed by Roger Penrose in 1979 [34]. This hypothesis was further supported by Marina Cortes and Lee Smolin [15], who introduced a large number of discrete states $S=\{I,J,K,L,\ldots\}$ forming energetic causal sets governed by an evolution rule. In this model, each state $I$ has a unique successor $J$. A similar framework is developed in the study of causal sets [7, 43], where discrete states $A,B,C,\ldots$ form a causal set $\mathcal{A}$, and the order $A\prec B\prec C\prec\ldots$ denotes a ”precedes” causal relation. This notion is widely discussed in the literature (see, e.g., [37]). The physical essence of a “state” can be characterized by energy and momentum [15], the discrete structure of 3D space [17], or entanglement entropy [42]. The fundamental role of entropy in this context was specifically investigated in [1].
@@ 3 The multidimensional arrow of time
The scale of branes is limited by the $D$-dim horizon. Such branes - solutions to classical equations (9) with the same action (8) describing the branes (18) and the expanding extra dimensions (11) have been studied separately, but with the same function of $f(R)$ gravity and the dimensionality of extra space, in [40] and [9]. The similar research can be found in [47], [21]. This approach is sufficient for our purpose, albeit at the cost of losing a smooth transition between the two regimes. The analogy follows the standard practice in cosmology: if our aim is to study the structure of a galaxy, we neglect processes like spatial expansion near the galaxy and solve the equations using the Minkowski metric. To study cosmological expansion, we consider the FRW metric as the solution to the Friedmann equations far from any local object. Matching these solutions at the boundary is of special interest and is not usually applied.
Let us address a reasonable question that arises at this stage. At first sight, the arrow of time appears to emerge from the global expansion far from the horizon. Conversely, an observer within the horizon cannot, by definition, perceive the total expansion. Why, then, is the arrow of time still recognized by such an observer?
To clarify this situation, we consider a similar problem within the framework of the McVittie metric, which describes a black hole of mass $m$ embedded in a $D$-dimensional de Sitter background. Its primary advantage for our purposes is that its line element is well-known and possesses a closed analytical form
$ds^{2}=\left(\frac{1-\frac{m}{4a(t)^{D-3}r^{D-3}}}{1+\frac{m}{4a(t)^{D-3}r^{D-3}}}\right)^{2}dt^{2}-a(t)^{2}\left(1+\frac{m}{4a(t)^{D-3}r^{D-3}}\right)^{\frac{4}{D-3}}(dr^{2}+r^{2}d\Omega^{2}_{D-2}),$ (19)
where $m$ represents the mass of the static compact object (e.g., a star or a black hole) and $a(t)=e^{Ht}$ is the scale factor of the background universe. To determine the region of gravitational dominance, we look for the radius $R_{c}$ where the inward gravitational pull exactly balances the outward de Sitter expansion. In $D$ dimensions, the physical radius of this ”static shell” is given by:
$R_{c}=\left(\frac{(D-3)m}{2H^{2}}\right)^{\frac{1}{D-1}}$ (20)
For any observer located at a physical distance $R(t)<R_{static}$, the geometric state of the manifold allows for traditional gravitational attraction toward the central mass $m$.
Therefore, in this single coordinate system, two distinct spatial regimes coexist. While the positive cosmological constant drives accelerated expansion on cosmological scales, its influence remains negligible in the strong-field region near the compact object. Consider an observer at rest at a physical radius $R$. If the observer is positioned below the critical radius, at $R<R_{crit}$, they are effectively decoupled from the global expansion. In this regime, the proper time $d\tau$ relates to the cosmological time $dt$ as:
$d\tau=\frac{1-\frac{m}{2r^{D-3}a(t)^{D-3}}}{1+\frac{m}{2r^{D-3}a(t)^{D-3}}}dt$ (21)
The factor in the r.h.s. is a monotonic function of both $t$ and $r$. Consequently, if $dt>0$, then $d\tau>0$; thus, the local observer experiences a well-defined proper arrow of time.
@@
In our model, the brane plays a role analogous to this decoupled region: the brane remains a stable, static entity residing deep “under the horizon,” while the global bulk continues its expansion. This configuration ensures that the established arrow of time is maintained on the brane despite its local stability.
@@
The Past Hypothesis postulates that the initial entropy of the Universe was extremely small compared to the entropy within the cosmological horizon today, [24]. Let us estimate their ratio within the framework of this work, using equation (17):
@@
Here, $S^{(D)}(0)$ denotes the Wald entropy contained within the cosmological horizon at the moment $t=0$ when our brane was created. $S^{(D)}(t)$ represents the Wald entropy of the same comoving volume at the present cosmic time, $t$.
@@
The entropy $S_{D}$, defined in Eq. (15), exhibits permanent growth in the multidimensional space. This constant flow of entropy dominates over any local entropy changes on the brane and supplies the arrow of time permanently. The initial value is extremely small compared to the present time and requires no fine-tuning within the framework of our model.
@@
A crucial feature of a multidimensional arrow is that it is robust. Even if our specific pocket universe were to undergo a ”Big Crunch” (where the local cosmological arrow might theoretically stall or appear to reverse), the multidimensional arrow remains unaffected. The global process of ”geometric nucleation” at scales much smaller than $10^{-27}$ cm ensures that the multiverse as a whole never reaches equilibrium.
@@ 5 Conclusion
This paper investigates the role of extra dimensions in the emergence of the arrow of time. As is well known, an initial volume, defined in a space with a $D$-dimensional de Sitter metric, grows continuously, generating disconnected volumes bounded by horizons. Each such volume is endowed with the Bekenstein-Hawking-Wald entropy. Consequently, the total entropy of the initial volume increases exponentially along the temporal coordinate.
@@
Observers are located on a static brane with a scale significantly smaller than the de Sitter horizon. Since matter is, by definition, confined to the brane, the bulk expansion remains unobservable to a local observer. However, the entropy growth of the extra dimensions provides a rigorous basis for the arrow of time. To this end, following Penrose’s idea, we introduce a Postulate of Causality related to the entropy of physical states. We demonstrate that the arrow of time arises from the fact that the multidimensional volume is a monotonic function of the temporal coordinate. Conversely, the observer’s proper time $\tau$ is a monotonic function of entropy (16); therefore, applying the Postulate of Causality establishes the arrow of time.
@@
Finally, let us distinguish between the cosmological arrow of time and the multidimensional one. The cosmological arrow of time originated during the inflationary stage within a horizon of $1/H\approx 10^{-27}$ cm and has since physically expanded to $1/H\approx 10^{28}$ cm at the present epoch. Its total entropy is stored in matter and, most significantly, within black holes.
In contrast, the multidimensional arrow of time is nucleated at much smaller scales, making its physical extent many orders of magnitude larger than the size of our observable universe. Its entropy, governed by the Bekenstein-Hawking-Wald framework, is stored within a constantly increasing number of pocket universes, each residing beneath its own causal horizon.
@@ Acknowledgment
The author is grateful to V. Berezin and A. Galiautdinov for their interest in the research. The work was partly funded by the Ministry of Science and Higher Education of the Russian Federation, Project ”Studying physical phenomena in the micro- and macro-world to develop future technologies”  FSWU-2026-0010 and the Kazan Federal University Strategic Academic Leadership Program.